OpenStudy (anonymous):

tan (3x-30(degrees)) = -1, 0

OpenStudy (anonymous):

if: \[\tan(\theta) = -1\] that means that: \[\theta=135,315\]degrees. So now we have: \[3x-30=135\iff 3x=165\iff x=55\]or\[3x-30=315\iff3x=345\iff x=115\]

OpenStudy (anonymous):

oh, ohhhh, got it. thanks :D